Allied Manufacturing has total assets of $4 million, financed with 50% debt and 50% equity. The cost of debt is 6% before taxes and the cost of equity capital is 11%. The company has EBIT of $300,000 and a tax rate of 30%. Estimate Allied’s residual income. (Enter your answer to the nearest $1,000. Leave the $ sign off. In other words, if your answer is $55,550, enter 56 for your answer. If your answer is negative, be sure to include a minus sign, like -56)
Answer #1

Interest to be paid = 4 million * 50% * 6% = 120000

EBT = 300000 - 120000 = 180000

Earnings after tax = 180000 * ( 1 - 0.30) = 126000

Required return on equity = 2000000 * 11% = 220000

Residual Income = 126000 - 220000 = - 94000